How to solve this sum using algebra equation.
A famer had 240 more ducks then chickens at first.He sold half of the ducks.After that he had 150 more chickens than ducks .How many ducks had he at first?

Answer by mananth(16949)

let the number of chickens he has be x
ducks = 240+x
(240+x) -1/2(240+x)
So he has 1/2 (240+x) ducks
1/2 (240+x) +150 = x
120+1/2 x+150= x
240 +x +300 /2 = x
540+x=2x
x=540 number of chickens
Number of ducks = 540+240=780